OpenStudy (anonymous):

A store had 175 cell phones in the month of January. Every month, 10% of the cell phones were sold and 10 new cell phones were stocked in the store. Which recursive function best represents the number of cell phones in the store f(n) after n months?

OpenStudy (anonymous):

(A) f(n) = 175 - 0.9 * f(n - 1) + 10, f(0) = 175, n > 0 (B) f(n) = 0.1 * f(n - 1) + 10, f(0) = 175, n > 0 (C) f(n) = 175 + 0.9 * f(n - 1) + 10, f(0) = 175, n > 0 (D) f(n) = 0.9 * f(n - 1) + 10, f(0) = 175, n > 0

OpenStudy (danjs):

so initially at time 0, n=0, f(n) = 175 phones

OpenStudy (danjs):

over a month 10% of them were sold, and 10 new ones were added 0.9*f(n-1) + 10 90% of the last value, plus 10

OpenStudy (danjs):

then when n increases to the next month, it will still be 90% of the last months still here , plust 10 new ones

OpenStudy (danjs):

recursive means that it will use the last value in the new equation, like, n=2 is used in figuring out the third month n=3 used to figure out the 4th month, all you need to know , is the initial value n=0, f(0)=175, and all the next months can be calculated

OpenStudy (danjs):

(D) f(n) = 0.9 * f(n - 1) + 10, f(0) = 175, n > 0

OpenStudy (anonymous):

Yeah D was what I was going with first as well.

OpenStudy (danjs):

if you put in n=1, the first month f(1) = 0.9* f(1-1) + 10 f(1) = 0.9*f(0) + 10 they say f(0) = 175 f(1) = 0.9*175 + 10

OpenStudy (danjs):

then that value for f(1), is used when you find f(2)

OpenStudy (danjs):

f(2) = 0.9*f(2-1) + 10 f(2) = 0.9* f(1) + 10 see the recursion?

OpenStudy (danjs):

f(3) will use the value for f(2) f(4) will use the value for f(3) . . .so on

OpenStudy (anonymous):

THANK YOU SO MUCH FOR HELPING ME UNDERSTAND!! :)

OpenStudy (danjs):

you are welcome, anytime :)